A television channel is assigned the frequency range from 54 MHz to 60 MHz. A series RLC tuning circuit in a TV receiver resonates in the middle of this frequency range. The circuit uses a 12 pF capacitor

A) In order to function properly, the current throughout the frequency range must be at least 50% of the current at the resonance frequency. What is the minimum possible value of the circuit's resistance?
Answer #1

Ric Criut R?+(Xx-XcJ? XL = 21FL. Xc= Eo Ic= R²+(La- At dreseuauce conddion, XL=XC Lwide cW. it (275)= LC LC 54 +60 2 57 X106* $=60M H. R= 271FL 2n%60Xn0 He )XO. 649X104)- _ ! 21tx(60X16X [12X103) R= 13.635e f=5#MH2- R. (2TTX (54X10S)X12XBey R=14.66